# Gamma-ray spectra of the Crab, Vela and Geminga pulsars fitted with SED of the emission from their current sheet

### 24 Jan 2024 v2

@@ 4.3 Characteristics of the source of the Geminga pulsar’s emission
for $\Delta\chi=10^{-3}$ (see Eq. (36) and recall that ${\hat{\kappa}}_{\rm obs}={\hat{\kappa}}_{\rm th}$). The dependence of ${\hat{B}}_{0}d^{2}$ on $\alpha$ is therefore given by a set of points whose only difference with those plotted in Fig. 5 is that their vertical coordinates are shifted upward by $0.065$ (cf. Eqs. (37) and (40)).
@@ 5 Concluding remarks
An up to date discussion of the emission mechanisms that have so far been proposed for explaining the data on the observed gamma-ray spectra of pulsars can be found in The H.E.S.S. Collaboration et al. (2023). Even in the most favourable circumstances, spectra of the emissions from relativistic charged particles that are accelerated by synchro-curvature, inverse Compton or magnetic reconnection processes fit the observed data only over disjoint limited ranges of photon energies. In particular, the constraints set by the $20$ TeV component of the radiation from the Vela pulsar “provide unprecedented challenges to the state-of-the-art models of HE [high energy] and VHE [very high energy] emission from pulsars” (The H.E.S.S. Collaboration et al., 2023).
In contrast, the SED of the tightly focused caustics that are generated by the superluminally moving current sheet in the magnetosphere of a non-aligned neutron star fits the observed spectra of the Crab, Vela and Geminga pulsars over the entire range of photon energies so far detected by Fermi-LAT, MAGIC and H.E.S.S. from them: over $10^{2}$ MeV to $20$ TeV. Not only are these data fitted by a single spectral distribution function but they are also explained by a single emission mechanism, an emission mechanism that has already accounted for a number of other salient features of the radiation received from pulsars: its brightness temperature, polarization, radio spectrum, and profile with microstructure and with a phase lag between the radio and gamma-ray peaks (Ardavan, 2021, 2022).
@@
It is often presumed that the plasma equations used in the numerical simulations of the magnetospheric structure of an oblique rotator should, at the same time, predict any radiation that the resulting structure would be capable of emitting (Spitkovsky, 2006; Kalapotharakos et al., 2012). This presumption stems from ignoring the role of boundary conditions in the solution of Maxwell’s equations. The boundary conditions with which the structure of the pulsar magnetosphere is computed are radically different from the boundary conditions with which the retarded solution of these equations (i.e. the solution describing the radiation from the charges and currents in the magnetosphere) is derived (see the last paragraph in Sect. 6 of Ardavan 2021).
Thickness of the current sheet, which sets an upper limit on the frequency of the present radiation, is dictated by microphysical processes that are not well understood: the standard Harris solution of the Vlasov-Maxwell equations (Harris, 1962) that is commonly used in analysing a current sheet is not applicable in the present case because the current sheet in the magnetosphere of a non-aligned neutron star moves faster than light and so has no rest frame. Even in stationary or subluminally moving cases, there is no consensus on whether equilibrium current sheets in realistic geometries have finite or zero thickness (Klimchuk et al., 2023). The fact that the SED described by Eq. (5) yields such good fits to the gamma-ray spectra of the pulsars analysed here lends support to treating the magnetospheric current sheet as volume-distributed but ultra-thin (see Ardavan, 2021, Sect. 4.7).
Acknowledgements.

### 03 Mar 2024 v3

@@ 4.3 Characteristics of the source of the Geminga pulsar’s emission
for $\Delta\chi=10^{-3}$ (see eq. (35) and recall that ${\hat{\kappa}}_{\rm obs}={\hat{\kappa}}_{\rm th}$). The dependence of ${\hat{B}}_{0}d^{2}$ on $\alpha$ is therefore given by a set of points whose only difference with those plotted in figure 5 is that their vertical coordinates are shifted upward by $0.065$ (cf. eqs. (36) and (39)).
@@ 5 Concluding remarks
Even in the most favourable circumstances, spectra of the emissions from relativistic charged particles that are accelerated by synchro-curvature, inverse Compton or magnetic reconnection processes fit the observed data only over disjoint limited ranges of photon energies. In particular, the constraints set by the $20$ TeV component of the radiation from the Vela pulsar “provide unprecedented challenges to the state-of-the-art models of HE [high energy] and VHE [very high energy] emission from pulsars” [2].
In contrast, the SED of the tightly focused caustics that are generated by the superluminally moving current sheet in the magnetosphere of a non-aligned neutron star fits the observed spectra of the Crab, Vela and Geminga pulsars over the entire range of photon energies so far detected by Fermi-LAT, MAGIC and HESS from them: over $10^{2}$ MeV to $20$ TeV. Not only are these data fitted by a single spectral distribution function but they are also explained by a single emission mechanism, an emission mechanism that has already accounted for a number of other salient features of the radiation received from pulsars: its brightness temperature, polarization, radio spectrum, profile with microstructure and with a phase lag between the radio and gamma-ray peaks [3, 11, 22] and the discrepancy between the energetic requirements of its radio and gamma-ray components [23].
@@
It is often presumed that the plasma equations used in the numerical simulations of the magnetospheric structure of an oblique rotator should, at the same time, predict any radiation that the resulting structure would be capable of emitting (e.g. [4, 5]). This presumption stems from disregarding the role of boundary conditions in the solution of Maxwell’s equations. As we have already pointed out, the far-field boundary conditions with which the structure of the pulsar magnetosphere is computed are radically different from the corresponding boundary conditions with which the retarded solution of these equations (i.e. the solution describing the radiation from the charges and currents in the magnetosphere) is derived (see section 3 and the last paragraph in section 6 of [3]).
Thickness of the current sheet, which sets a lower limit on the wavelength of the present radiation, is dictated by microphysical processes that are not well understood: the standard Harris solution of the Vlasov-Maxwell equations [24] that is commonly used in analysing a current sheet is not applicable in the present case because the current sheet in the magnetosphere of a non-aligned neutron star moves faster than light and so has no rest frame. Even in stationary or subluminally moving cases, there is no consensus on whether equilibrium current sheets in realistic geometries have finite or zero thickness [25]. The fact that the SED described by eq. (5) yields such good fits to the observed spectra of the pulsars analysed here corroborates the notion that, though necessarily volume-distributed [10], the magnetospheric current sheet is sufficiently thin to generate gamma rays (see section 4.7 of [3]).
acknowledgement